If there are 0.0345 mols of HI in 47 mL of solution, what is its concentration? Also, what
is the pH?
1
Expert's answer
2020-05-03T15:24:05-0400

C = n/V

C = 0.0345/0.047 = 0.734 M

HI is a strong acid and dissociates completely.

HI -> H+ + I-

[H+] = 0.734 M

pH = -Log[H+] = -Log(0.734) = 0.134